A 12oz can of soft drink (assume m= 340 g) at 25 degrees Celsius is placed in a freezer where the temperature is -12 degrees Celsius. How much energy must be removed from the soft drink for it to reach this temperature?
1
Expert's answer
2019-10-09T07:36:57-0400

"q=cm(T_2-T_1)"


We should make an assumption that specific heat of soft drink is the same as that of water: "c=4.186 \frac{J}{g\times \circ C}"


"q=4.186\times340\times(-12-25)= -52.7 kJ"